# Joint Denoising and Fusion with Short- and Long-exposure Raw Pairs

### v4

@@ Abstract
Denoising and high dynamic range (HDR) imaging are significant yet challenging problems due to the small aperture and sensor size of generic image sensors. Current methods predominantly generate HDR images from a set of bracketed exposure sRGB images. However, they overlook the computational and memory inefficiencies of the Image Signal Processor (ISP) when processing a set of sRGB images with different exposures. Furthermore, the absence of large-scale raw-based HDR datasets limits the research on HDR imaging. Unlike existing methods, the core idea of this work is to utilize the difference between short- and long-exposure images of signal-to-noise ratios to generate HDR images and denoising. To this end, we propose a model tailor-made for double-exposure HDR sensors, leveraging the unique features of the raw data to facilitate raw-to-HDR mapping and raw denoising. Our key insights are threefold: (1) a new computational raw LDR-HDR pair formation pipeline is designed to construct a real-world raw HDR dataset called RealRaw-HDR; (2) a lightweight-efficient HDR model, RepUNet, is developed using the structural reparameterization technique; (3) a plug-and-play alignment-free and motion-aware short-exposure-first selection loss and a colorfulness loss are proposed to mitigate ghost artifacts and color cast. Our empirical evaluation validates the effectiveness of the proposed LDR-HDR formation pipeline, as well as experiments show that our method achieves comparable performance to the state-of-the-art methods with less computational cost.
@@ 1 Introduction
Figure 1: To address the HDR reconstruction problem, previous methods mainly focus on models trained on processed sRGB datasets, while we promote a novel view to address HDR reconstruction by training on raw images. Furthermore, compared with sRGB-based HDR reconstruction, raw-based HDR reconstruction is more beneficial and reduces the power consumption of ISP, DDR, and NPU by almost half.
@@
Most resource-constrained cameras exhibit low dynamic ranges (LDR), rendering them unable to capture the full range of brightness and color information in real-world scenes. Conversely, high dynamic range imaging seeks to encompass a significantly broader range of luminance values, compensating for color distortions and the subtle detail loss observed in LDR images. Despite the availability of dedicated hardware for directly acquiring HDR images, such equipment is typically expensive, thus limiting its practicality for most users. As a result, there has been an increasing focus on fusion-based HDR imaging methods tailored for LDR cameras.
Widely practiced HDR reconstruction methods [16, 5, 27, 37, 39] exploit a set of sRGB images with different exposure levels to produce an HDR image. However, capturing a set of bracketed exposure LDR images can be challenging, especially when the camera is shaking or there are moving objects in the viewfinder frame. [16]. In response to this challenge, Sony has introduced a Digital Overlap (DOL) [41] HDR sensor and a Binned Multiplexed Exposure (BME) HDR sensor [24]. These sensors can shorten the time interval between different exposure images, enabling the “quasi-synchronous” capture of multiple raw images with different exposure times. Nonetheless, the quasi-synchronous mechanism retains relative motion between pre- and post-frames, which can introduce ghost artifacts in the reconstructed HDR results.
Although sRGB-based HDR reconstruction methods [38, 39] have made significant progress in recent years, these methods neglect three critical considerations. Firstly, as shown in Fig. 1, an ISP pipeline needs to be applied separately to different exposure LDR raw images to obtain different exposure LDR sRGB images, which are then reconstructed as an HDR image. The ISP pipeline typically involves complicated processes, including demosaicing, tone mapping, color and contrast tuning, denoising and sharpness, etc. These processes require significant memory and computational resources, which is very unfriendly to resource-limited sensors like security cameras. Secondly, existing approaches [16, 21] substitute the complex camera response curve with a simplistic inverse gamma function to obtain a set of pseudo-linear sRGB images with different exposures. However, this approach may produce non-luminance-aligned images in non-saturated regions during luminance alignment, thus rendering training more challenging. The linear characterization between the different exposure images of the raw images reduces the difficulty of luminance-aligned processing. Thirdly, the raw image data contains more detailed information, and the noise model is more straightforward. Meanwhile, the bit depth of the raw image is generally higher than that of the sRGB image, which is more beneficial to improving the dynamic range of the image. Therefore, the limitations inherent in the current sRGB-based HDR reconstruction methods highlight the necessity of further research and development in the domain of HDR imaging.
Building upon the above observations, we propose reconstructing HDR images from a set of raw LDR images with different exposures (as shown in Fig. 1. Accordingly, we introduce a novel raw LDR-HDR pair formation pipeline and obtain an associated training and testing dataset named RealRaw-HDR. In addition, we devise a simple yet effective raw HDR reconstruction network with reparameterization techniques, establishing a new baseline for the raw HDR reconstruction field. Specifically, upon exploring the HDR sensor imaging system, we discover that simulating the sensor’s Digital Gain can produce the same luminance level as altering the exposure time, given certain conditions. Leveraging this insight, we design a digital imaging pipeline that captures raw images with changing exposure gain based on the camera’s photoelectric signal conversion model. By simulating the HDR sensor imaging system, we built a digital imaging pipeline to obtain raw images with different exposures, thus emerging as a practical tool for generating high-quality LDR-HDR images. The resultant RealRaw-HDR dataset encompasses a substantial collection of LDR-HDR pairs, serving as the foundation for training and evaluation. By incorporating the distinctive attributes of raw images into our approach, we can achieve elevated HDR reconstruction outcomes characterized by enhanced efficiency and accuracy.
@@
To further advance the HDR imaging system, we introduce RepUNet, a lightweight and resource-efficient model for raw HDR reconstruction. RepUNet adopts reparameterization techniques, bypassing the need for computationally intensive alignment modules such as optical flow, deformation convolution, or attention mechanisms, which are commonly employed in existing deep learning-based HDR reconstruction methods [16, 21]. Furthermore, addressing the absence of explicit alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss (AMSS-Loss), which guides the network to focus on local motion regions, effectively mitigating ghost artifacts between short- and long-exposure images. This approach significantly reduces hardware expenses and enhances the real-time performance of HDR imaging systems.
@@
    We propose a pioneering pipeline for forming LDR-HDR pairs based on sensor imaging theory, resulting in a large-scale, high-quality raw HDR dataset coined RealRaw-HDR. This dataset serves as the bedrock for follow-up raw HDR reconstruction methodologies.
@@
    We develop a lightweight model, RepUNet, to achieve raw HDR reconstruction, which utilizes structural reparameterization techniques that notably curb the computational demands of hardware devices.
@@
    We introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate ghost artifacts between short- and long-exposure images.
Our contributions signify a substantial advancement in raw HDR image reconstruction research, delivering an efficient and productive solution for producing superior HDR images. Extensive experiments demonstrate that our approach achieves state-of-the-art visual quality and quantitative metrics performance.
@@ 2.1 HDR Imaging
The expansion of camera dynamic range has been a focal point of research over the past decades. One class of technology [33, 31] involves achieving HDR image capture in a single shot by refining digital camera hardware design. The Dual Conversion Gain (DCG) sensor [15], for instance, enhances dynamic range through DCG circuitry, enabling image capture in extremely low-light conditions without compromising performance in high-light scenarios. Sun et al. [33] propose a method for single-exposure HDR imaging through end-to-end optical coding design. Regrettably, these methods often necessitate customization to specific optical systems or image sensors. HDR techniques relying on hardware modification are typically expensive and beyond the reach of average users.
Another class of approaches [25, 21, 36] is time-domain multi-frame fusion HDR, which merges a set of bracketed exposure LDR sRGB images to produce an HDR image. However, this process—capturing images from underexposed to overexposed configurations can result in blurring artifacts for moving objects. Sony has endeavored to address this by introducing the Digital Overlap (DOL) HDR sensor [6]. For instance, dual-exposure DOL sensors like Sony IMX327, IMX385, and OV OS05B produce long-exposure and short-exposure frames in line-interleaving patterns. This pattern significantly shortens the interval between these frames, thus minimizing relative motion. Presently, many methods use DOL sensors to capture a set of bracketed exposure raw images, which ISP processes, and subsequently fusion bracketed exposure LDR sRGB images to reconstruct an HDR image.
Recent trends in HDR imaging involve the application of deep neural networks [18, 26, 2] to learn the production of high-quality HDR images from a set of bracketed exposure LDR images. Kalantari et al. [16] propose a CNN-based HDR approach that employs optical flow to align LDR images before network inference. Wu et al. [36] approach HDR imaging as an image translation problem without explicit motion alignment. Yan et al. [37] introduce spatial attention to achieve LDR image alignment. Liu et al. [21] present an attention-guided deformable convolutional network for multi-frame HDR imaging. Prabhakar et al. [29] introduce an efficient method for generating HDR images using a bilateral guided up-sampler and explore zero-learning for HDR reconstruction. Niu et al. [27] propose a multi-frame HDR imaging method based on generative adversarial learning. Liu et al. [22] propose a Transformer-based [28] HDR imaging method. These deep learning-based approaches [37, 21, 22] consistently elevate state-of-the-art performance. However, these methods reconstruct HDR based on sRGB images at the end of the ISP pipeline and only train on one dataset. They overlook the large computational and storage resources required by the ISP pipeline to process bracketed exposure raw images. It also complicates the ISP system, making it challenging for resource-limited cameras to output high-quality video/images.
@@ 2.2 Low-level Raw Image Processing
Learning-based processing of low-level raw images [19, 43] has received increasing attention. Google’s HDR+ [11] produces the raw HDR image by aligning and merging a burst of raw frames with the same low exposure. Zhang et al. [44] construct a real-world super-resolution dataset by designing an optical zoom system and propose a baseline network with a bilateral contextual loss. Qian et al. [30] solve the joint demosaicing, denoising, and super-resolution task with the raw input. Wang et al. [34] propose a lightweight and efficient network for raw image denoising. Sharif et al. [1] offer a new learning-based approach to tackle the challenge of joint demosaicing and denoising on image sensors. Wei et al. [35] investigate the low-light image denoising considering the sensor’s photoelectric properties. Yue et al. [42] achieve state-of-the-art raw image denoising by constructing a dynamic video dataset with noise-clean pairs.
Learning-based raw image processing has demonstrated outstanding potential for high-performance reconstruction from raw sensor data. However, acquiring paired data in the raw domain is difficult and expensive. Our work proposes a new large-scale, high-quality raw dataset and provides a pipeline to acquire raw LDR-HDR pairs based on the imaging system.
@@ 3 New raw LDR-HDR pair Formation Pipeline
We begin by analyzing the sensor response of the imaging system and presenting a novel data formation pipeline for raw LDR-HDR pairs based on the camera response model.
@@ 3.1 Analysis of CMOS Imaging System
Given that $(x,y)$ in Eq. 1 pertains to a single photosensory cell, we assume that each parameter remains constant with respect to position. Consequently, the coordinates $(x,y)$ can be omitted [14]:
@@
Subsequently, the camera’s amplifier circuit amplifies the electrical signal, yielding the raw camera response value through analog-to-digital conversion [47]:
@@
We observe from Eq. 5 that under noise-free conditions, adjusting the gain factor ($K_{\textrm{a}},K_{\textrm{d}}$) can linearly adjust the camera’s raw response value. This linear characteristic allows us to achieve an equivalent result as modifying the exposure time ($T$) by simulating the Digital Gain, thereby obtaining a bracketed exposure raw image.
Nonetheless, due to noise in real-world imaging, changing the Digital Gain will also linearly amplify the noise. Therefore, to mitigate noise influence, we avoid this problem as much as possible during data acquisition (in Sec. 3.2.1). Also, we avoid using multiplicative gain to construct long-exposure images that amplify the noise difference between long- and short-exposure images and instead use truncation to similar long-exposure images. In addition, we synthesize the long-short exposure images by measuring the accurate HDR sensor noise model to eliminate the long-short exposure image noise difference. Based on this rationale, we propose an innovative and efficient pipeline for generating LDR-HDR data pairs.
3.2 Formation of Raw LDR-HDR Image Pair
Relative to sRGB images, HDR reconstruction from raw images offers several distinct advantages: it retains a more significant amount of original information and has higher bit depths, which is more beneficial for improving the dynamic range of the image; the Image Signal Processing (ISP) pipeline is more straightforward and less computational demands. These characteristics render raw HDR reconstruction a highly promising approach for deployment in edge devices. Considering these considerations, we create a raw HDR dataset featuring a collection of LDR-HDR data pairs, aptly named RealRaw-HDR.
@@ 3.2.1 Data Acquisition
Based on the analysis in Sec. 3.1, we observe that changing the Digital Gain yields a comparable luminance to adjusting the exposure time when the noise level remains sufficiently low. Consequently, we can capture two raw images with identical exposure settings and subsequently manually simulate the Digital Gain to generate two raw images with differing exposures. It’s crucial to emphasize that these two raw images with the same exposure require low noise and no overexposure. In order to capture near-noise-free raw images, we utilize high-end cameras, special photography techniques, and post-processing acquisitions.
Specifically, we use a top-of-the-line FUJI-FILM GFX50S II camera with a wide-aperture lens to capture high-quality raw images. The camera has a $51$ megapixel medium format image sensor with a pixel size of $5.3\mu m$ (The iPhone 14 Pro Max primary camera single pixel size is only $1.22\mu m$). We also set the camera’s ISO to 800 or below and turned on the noise reduction feature to enhance image quality. Subsequently, we subject the raw image to post-processing using the BM3D algorithm, attaining nearly noise-free raw images.
Meanwhile, to simulate the relative motion between long and short exposure images within a real-world scene, we use a human subject to simulate motion between LDR images and trigger the shutter twice in a rapid time interval. To eliminate the risk of unintended camera shake, we securely mount the camera on a tripod and use a remote smartphone to control the shutter release. Following image capture, the raw images undergo black level correction and normalization before being subjected to BM3D processing [23] to effectively reduce noise, thereby yielding raw images nearly devoid of noise.

### v5

# Efficient HDR Reconstruction from Real-World Raw Images

@@ Abstract
The widespread usage of high-definition screens on edge devices stimulates a strong demand for efficient high dynamic range (HDR) algorithms. However, many existing HDR methods either deliver unsatisfactory results or consume too much computational and memory resources, hindering their application to high-resolution images (usually with more than 12 megapixels) in practice. In addition, existing HDR dataset collection methods often are labor-intensive. In this work, in a new aspect, we discover an excellent opportunity for HDR reconstructing directly from raw images and investigating novel neural network structures that benefit the deployment of mobile devices. Our key insights are threefold: (1) we develop a lightweight-efficient HDR model, RepUNet, using the structural re-parameterization technique to achieve fast and robust HDR; (2) we design a new computational raw HDR data formation pipeline and construct a real-world raw HDR dataset, RealRaw-HDR; (3) we propose a plug-and-play motion alignment loss to mitigate motion ghosting under limited bandwidth conditions. Our model contains less than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU. While being highly efficient, our model also outperforms the state-of-the-art HDR methods in terms of PSNR, SSIM, and a color difference metric.
@@ 1 Introduction
Most resource-constrained cameras exhibit low dynamic ranges (LDR), rendering them unable to capture the full range of brightness and color information in real-world scenes. Conversely, high dynamic range (HDR) imaging seeks to encompass a significantly broader range of luminance values, compensating for color distortions and the subtle detail loss observed in LDR images. Despite the availability of dedicated hardware for directly acquiring HDR images, such equipment is typically expensive, thus limiting its practicality for most users. As a result, there has been an increasing focus on fusion-based HDR imaging methods.
Recent methods [35, 20, 4, 25, 34] based on convolution neural networks (CNNs) [17, 22] have made impressive progress in HDR reconstruction performance, thanks to their scalability and flexibility from constructing elementary building blocks like convolutional layers. However, superior performance is usually obtained at a cost of heavy computational burden [20, 34, 21]. Although this can be alleviated by elaborate network structures or dedicated computing engines (e.g., GPU and NPU), the hardware cost and power consumption still limit the deployment of existing deep HDR reconstruction networks. Specifically, the growing number of high-definition screens on edge devices (e.g., smartphones, security cameras, and televisions) calls for a practical HDR reconstruction solution.
On the other hand, in the image processing pipeline, HDR reconstruction is widely used in the sRGB domain. Previous methods [35, 22, 36] exploit a set of sRGB images with different exposure levels to produce an HDR image, which has made rapid development in recent years. However, they tend to overlook three critical aspects. 1) Dataset Collection: Existing methods [3, 46] follow Kalantari et al. [14] to construct datasets. They first make the subject static and take three sets of images with different exposures, and then make the subject move twice to take dynamic images with different exposures. However, this process is labor intensive and difficult to acquire on a large scale. 2) ISP Processing Speed: When obtaining raw LDR images with different exposures, the ISP pipeline must be performed separately on each exposure. This incurs additional memory and computational overhead and leads to lower frame rates for HDR image output. 3) Reconstruction Quality: Raw images contain more delicate details of the original sensor signal that can be lost while processing sRGB images. The limitations of current HDR reconstruction methods highlight the need for further research and development in HDR imaging.
@@
In this paper, we propose an efficient scheme for HDR image reconstruction in the raw image domain. By analyzing the HDR image sensor system, we design a lightweight and efficient model for raw HDR reconstruction named RepUNet. RepUNet adopts reparameterization techniques and does not contain explicit computationally expensive alignment modules, such as optical flow [14], deformation convolution [4], or attention [20, 37], which are commonly used in existing deep learning-based HDR reconstruction methods [14, 20, 33, 24]. To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss, which encourages the network to focus on local motion patterns and alleviate misalignment between short- and long-exposure images. Consequently, our approach significantly reduces hardware costs and improves the real-time performance of HDR imaging systems.
To further promote the HDR imaging system, we investigate the HDR sensor imaging principle. We observe that changing the Gain of the image sensor can have a similar effect as modifying the exposure time under noise-free conditions. Leveraging this insight, we design an automatic control imaging system that captures raw images with different exposures, based on a digital camera photoelectric signal conversion model. This automatic control operable system satisfies real-world scenes’ dynamic range requirements, making it a practical tool for generating high-quality HDR images. The resulting RealRaw-HDR dataset includes many LDR-HDR pairs for training and evaluation. By incorporating the unique characteristics of raw images into our approach, we can achieve superior HDR reconstruction results with increased efficiency and accuracy.
@@
    We investigate the structure re-parameterizable technique for the HDR task and propose a lightweight model, RepUNet, with Topological Convolution Block (TCB). TCB can be used to improve the HDR performance of any HDR model without introducing any extra burden for inference.
@@
    We introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate ghost artifacts.
@@
    We propose a novel computational photography-based pipeline for raw HDR image formation and construct a real-world raw HDR dataset, i.e., RealRaw-HDR.
Our contributions represent a significant step forward in raw HDR image reconstruction research, providing an effective and efficient solution for producing high-quality HDR images. Our model contains less than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU. While highly efficient, our model also outperforms the state-of-the-art HDR methods by a large margin in terms of PSNR, SSIM, and a color difference metric.
@@ 2.1 HDR Imaging
Recently, benefiting from the fast development of deep learning techniques, training deep neural networks for effective HDR reconstruction has become increasingly popular. Many methods apply deep neural networks [16, 24, 29, 13] to learn the production of high-quality HDR images from a set of bracketed exposure LDR images. Kalantari et al. [14] proposed a CNN-based HDR approach that employs optical flow to align LDR sRGB images before network inference. Wu et al. [33] approached HDR imaging as an image translation problem without explicit motion alignment. Yan et al. [34] introduced spatial attention to achieve LDR image alignment. Liu et al. [20] presented an attention-guided deformable convolutional network for multi-frame HDR imaging. Prabhakar et al. [27] introduced an efficient method for generating HDR images using a bilateral guided up-sampler and exploring zero-learning for HDR reconstruction. Niu et al. [25] proposed a multi-frame HDR imaging method based on generative adversarial learning. Liu et al. [21] proposed a Transformer-based [26] HDR imaging method. These deep learning-based approaches [34, 20, 21] consistently elevated state-of-the-art performance. However, these methods reconstruct HDR based on sRGB images at the end of the ISP pipeline and only train on one dataset. They overlook the large computational and storage resources the ISP pipeline requires to process bracketed exposure raw images. It also complicates the ISP system, making it challenging for resource-limited cameras to output high-quality video/images.
To simplify the ISP system, another class of HDR reconstruction methods is based on raw image input. Google HDR+ produced the raw HDR image by aligning and merging a burst of raw frames with the same low exposure. Nevertheless, this approach requires a complex ISP system design and takes up a lot of DDR memory. Zou et al. [46] proposed reconstructing HDR images from a single raw image and collecting a raw/HDR paired dataset. However, this dataset is not suitable for real HDR sensors. Therefore, none of the existing methods can meet the requirements of real scenarios.
@@ 2.2 Low-level Raw Image Processing
Due to the merits of raw data, raw-based image processing [18, 41] has made significant progress in recent years. The work in [40] first performs the demoireing task in the raw domain and then utilizes a pretrained ISP module to transform the result into the sRGB domain. Yang et al. [38] proposed a single-stage network empowered by feature domain adaptation to decouple the denoising and color mapping tasks in raw low-light enhancement. Zhang et al. [42] constructed a real-world super-resolution dataset by designing an optical zoom system and proposed a baseline network with a bilateral contextual loss. Qian et al. [28] solved the joint demosaicing, denoising, and super-resolution task with the raw input. Wang et al. [30] proposed a lightweight and efficient network for raw image denoising. Sharif et al. [1] proposed a new learning-based approach to tackle the challenge of joint demosaicing and denoising on image sensors. Wei et al. [31] investigated the low-light image denoising considering the sensor photoelectric properties. Yue et al. [39] achieved state-of-the-art raw image denoising by constructing a dynamic video dataset with noise-clean pairs. Learning-based raw image processing has demonstrated outstanding potential for high-performance reconstruction from raw sensor data. However, acquiring paired data in the raw domain is difficult and expensive. Our work proposes a new large-scale, high-quality raw dataset and provides a pipeline to acquire raw LDR-HDR pairs based on the imaging system.
@@ 3 New raw LDR-HDR pair Formation Pipeline
We first analyze the sensor response of the imaging system and propose a new formation pipeline for raw HDR-paired data based on the camera response model.
@@ 3.1 Analysis of CMOS Imaging System
Given that $(x,y)$ in Eq. 1 pertains to a single photosensory cell, we assume that each parameter remains constant concerning position. Consequently, the coordinates $(x,y)$ can be omitted [12]:
@@
Subsequently, the camera amplifier circuit amplifies the electrical signal, yielding the raw camera response value through analog-to-digital conversion [45]:
@@
We observe from Eq. 5 that under noise-free conditions, adjusting the gain factor ($K_{\textrm{a}},K_{\textrm{d}}$) can linearly change the camera raw response value. This linear characteristic allows us to achieve an equivalent result to modifying the exposure time ($T$) by simulating the gain, thereby obtaining a set of bracketed exposure raw images. However, there is unavoidable noise in the actual imaging process. Therefore, we follow the existing denoising methods [39, 32] and try to avoid the effect of noise as much as possible during data acquisition (in Sec. 3.2.1).
3.2 Formation of Short- and Long-exposure Raw Pairs
Compared to sRGB images, HDR reconstruction from raw images has the advantages of more original information, simpler ISP processing, and less computation, making it a promising paradigm to deploy in edge devices. To this end, we construct a new raw HDR dataset with LDR-HDR data pairs, named RealRaw-HDR.
@@ 3.2.1 Data Acquisition
Based on the analysis in Sec. 3.1, we find that changing the sensor digital Gain, $K_{\textrm{d}}$, can achieve a similar luminance to adjusting the exposure time $K_{\textrm{a}}$ on the noise-free condition. Consequently, we use a top-of-the-line FUJI-FILM GFX50S II camera with a wide-aperture lens to capture high-quality raw images. The camera has 15.5 stops of dynamic range (15.5 bit) and a $51$ megapixel medium format image sensor with a pixel size of $5.3\mu m$ (The iPhone 15 Pro Max primary camera single pixel size is only $1.22\mu m$). We also set the camera ISO to 800 or below and turned on the noise reduction feature to enhance image quality. At this point, the captured raw image has a low noise level.
Specifically, we capture two raw images ($I_{1}$ and $I_{2}$) with the same exposure settings using a high-end camera. Meanwhile, we use a human subject to simulate motion between images and trigger the shutter twice in a rapid time interval, to simulate the relative motion between short- and long-exposure images within a dual-exposure sensor. Further, to eliminate the risk of unintended camera shake, we mount the camera on a tripod and use a remote smartphone to control the shutter release. Afterward, the raw images are black level corrected, normalized, and then processed with BM3D [23] to reduce noise, which obtains nearly noise-free raw images. Note that there is a small relative motion between these two raw images, which is common in multi-frame HDR reconstruction. Our dataset will be released after the acceptance of this work.
Figure 1: The raw LDR-HDR pair formation pipeline. Two clean high dynamic range raw images, $I_{1}$ and $I_{2}$, have been processed through black-level correction, and normalization. After manual digital gain, clip, add noise, and normalization, the long-exposure image $I_{l}$ is overexposed in the bright areas, and the short-exposure image $I_{s}$ dark area information is covered by noise.